Find the equation of the median of triangle ABC through B if A= (2,4) B= (2,-3) C= (-4,-2) are the vertices.

I tried answering it and got the equation as 2x-3y=8=0 .

Is this the right answer?

Grade:11

3 Answers

Aman Bansal
13 years ago

DEAR SUHAIL,

MEDIAN IS A LINE SEGMENT PASSING THROUGH A VERTEX AND TO THE MIS POINT OF OPPOSITE LINE

THEREFORE THE EQUATION MUST PASS THROUGH B(2,-3)

AND ALSO MID POINT OF LINE SEGMENT BETWEEN POINTS (2,4) AND (-4,-2) THAT IS (-2, 1)

IT IS CLEAR THAT YOUR EQUATION DOES NOT SATISFY THE POINTS

THE CORRECT EQUATION CAN BE CALCULATED USING TWO POINT FORM PASSING THROUGH (2,-3)

AND (-2,1)

THE EQUATION IS x+y=-1

BEST OF LUCK

We are all IITians and here to help you in your IIT JEE preparation.
Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.
Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar : Click here to download the toolbar..  
AMAN BANSAL

suhail b
13 years ago

I did not understand how did you get that equation using two point formula???

xyz xz
12 years ago

Median: line passing through vertex and mnidpoint of opposite side

Eqn of median through A is 13x-6y=2

Eqn of median through B is 4x+3y+1=0

Eqn of median through C is 5x-12y=4